Stan has a $18,200 certificate of deposit that earns 4.8% interest compounded quarterly for 5 years. What is the compound interest after 5 years?
Answer by Theo(13342)

formula for discrete compounding is f = p * (1 + r) ^ n

f is the future value
p is the present value
r is the interest rate per time period
n is the number of time periods.

your time periods are in quarters.

p = 18,200
r = 4.8% / 4 = 1.2% per month / 100 = .012 per quarter.
the formula requires interest rate, not interest rate percent.
interest rate percent / 100 = interest rate.
n = 5 years * 4 = 20 quarters

formula becomes f = 18200 * (1 + .012) ^ 20 = 23103.70539533.

the interest earned is equal to f - p = 23103.70539533 - 18,200 = 4903.70539533.

that's your solution.